OpenStudy (anonymous):

PLEASE HELP!! Find the rectangular coordinates of the point with the polar coordinates [7, 2π/3] 1) [-7/2, 7√3/2] 2) [-7/2, 7/2] 3) [7/2, -7/2] 4) [7√3/2, -7/2]

zepdrix (zepdrix):

\[\Large\rm (r,\theta)=\left(7,\frac{2\pi}{3}\right)\] \[\Large\rm x=r \cos \theta,\qquad\qquad y=r \sin \theta\]

zepdrix (zepdrix):

\[\Large\rm x=7\cos\frac{2\pi}{3}\]Remember your special angles and the unit circle and all that?

OpenStudy (anonymous):

A little bit.